# Intuitive Fine-Tuning: Towards Simplifying Alignment into a Single Process

### v2

@@ Abstract
Supervised Fine-Tuning (SFT) and Preference Optimization (PO) are two fundamental processes for enhancing the capabilities of Language Models (LMs) post pre-training, aligning them better with human preferences. Although SFT advances in training efficiency, PO delivers better alignment, thus they are often combined. However, common practices simply apply them sequentially without integrating their optimization objectives, ignoring the opportunities to bridge their paradigm gap and take the strengths from both. To obtain a unified understanding, we interpret SFT and PO with two sub-processes — Preference Estimation and Transition Optimization — defined at token level within the Markov Decision Process (MDP) framework. This modeling shows that SFT is only a specialized case of PO with inferior estimation and optimization. PO evaluates the quality of model’s entire generated answer, whereas SFT only scores predicted tokens based on preceding tokens from target answers. Therefore, SFT overestimates the ability of model, leading to inferior optimization. Building on this view, we introduce Intuitive Fine-Tuning (IFT) to integrate SFT and Preference Optimization into a single process. IFT captures LMs’ intuitive sense of the entire answers through a temporal residual connection, but it solely relies on a single policy and the same volume of non-preference-labeled data as SFT. Our experiments show that IFT performs comparably or even superiorly to sequential recipes of SFT and some typical Preference Optimization methods across several tasks, particularly those requires generation, reasoning, and fact-following abilities. An explainable Frozen Lake game further validates the effectiveness of IFT for getting competitive policy. Our code will be released at https://github.com/TsinghuaC3I/Intuitive-Fine-Tuning.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable powerful potential across various downstream tasks after pre-training on large-scale corpora brown2020language ; achiam2023gpt ; zhou2024generative . However, their instruction-following skills and trustworthiness still fall short of expectations bender2021dangers ; bommasani2021opportunities ; li2022trust . Therefore, algorithms such as Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) ziegler2019fine ; ouyang2022training ; lee2023rlaif are used to further enhance LLMs’ abilities and align them better with human preferences.
@@
Figure 1: Comparison of Alignment Methods. IFT conducts alignment solely relying on positive samples and a single policy, starting from a pre-trained base model. IFT not only shows high efficiency as SFT, but also performs comparably or even superiorly to other alignment algorithms.
@@
Recently, Direct Preference Optimization (DPO) rafailov2024direct was proposed to integrate Reward Modeling and Policy Optimization into one single procedure using a loss function derived from Proximal Policy Optimization (PPO) schulman2017proximal . This approach demonstrates the potential to unify SFT and RLHF for the first time. Henceforth, many extended methods have been tried to realize this objective by bridging the gap between SFT and DPO. Some of them ethayarajh2024kto ; hong2024reference ; zhang2024negative aim to transform the contrastive loss of DPO into a SFT-like cross-entropy loss, learning positive samples similar to SFT while unlearning negative samples resort to Unlikelihood Training welleck2019neural . Some others get rid of the preference-labeling process before training, switching to collect samples and labels/rewards in an online manner liu2023statistical ; yuan2024self ; guo2024direct ; calandriello2024human ; tajwar2024preference , or just treating the SFT targets and online policy generations as positive and negative samples respectively xiong2023iterative ; chen2024self ; mitra2024orca ; liu2024extensive . Nevertheless, preference-labeled pairwise data is still essential in these methods, and the need for reference model only becomes unnecessary in some cases. Thus the core differences between SFT and Preference Optimization are not eliminated thoroughly. To address this challenging issue, a deeper and more unified understanding of them are needed.
@@
In this paper, we attempt to explain the similarities and differences between SFT and Preference Optimization by defining Preference Estimation and Transition Optimization in terms of state-action pairs within the Markov Decision Process (MDP) framework. Through this modeling, we demonstrate that SFT is simply a specialized case with inferior estimation and optimization among all Preference Optimization methods. To be specific, typical Preference Optimization methods usually sample the entire answer of policy based solely on the initial instruction, and align the policy preference reflected by the sampled sentence with human preferences. In contrast, SFT only samples a single token based on intermediate state of target answer, which leads to a biased estimation of policy preference and an inferior alignment shortcut. Or put it in a metaphorical view, when preparing for an exam, it is obviously more beneficial to review reference answers after completing entire questions, rather than trying to infer each subsequent step based on the reference answer ahead and checking it immediately. It is the similar case in LMs’ training, where SFT is less beneficial for enhancing LMs.
@@
Depending on this understanding, we introduce a unified alignment algorithm named Intuitive Fine-Tuning (IFT), which builds policy preference estimation in an intuitive manner like human, who can have a vague sense of the complete answer after hearing a question. Through a closer estimation to truly policy preference than SFT, IFT achieves a comparable or even superior alignment performance compared to the sequential recipe of SFT and Preference Optimization. Additionally, IFT requires only a single policy model, and the same volume and format of data as SFT, enjoying both data and computation efficiency. These characteristics also enable IFT to be implemented in domains where preference data is unavailable or expensive to collect.
Our main contribution are three folds: 1) By modeling within the MDP framework, we explain the similarities and differences between SFT and two fundamental Preference Optimization methods, PPO and DPO; 2) Based on this modeling, we introduce Intuitive Fine-tuning (IFT), a deeply unified version of SFT and Preference Optimization. IFT utilizes temporary residual connections to extract the model’s generation preference given the initial instructions, offering similar efficiency to SFT while achieving performance closer to Preference Optimization; 3) Through experiments on several benchmarks, we validate that IFT performs comparably or even superiorly to various SFT recipes and existing Preference Optimization methods. In particular, IFT demonstrates significantly stronger reasoning and fact-following abilities than the chosen baselines. An explainable toy-setting Frozen Lake further demonstrates the effectiveness of IFT.
2 Related Work
Classical Reinforcement learning (RL) has demonstrated strong performance in various sequential decision-making and optimal control domains, including robotics levine2018learning , computer games vinyals2019grandmaster and others guan2021direct . There are two main categories of RL algorithms: value-based and policy-based, depending on whether they learn a parameterized policy. Value-based RL aims to fit an value function defined by Bellman Equation, containing methods such as Monte-Carlo (MC) Learning lazaric2007reinforcement and Temporal Difference Learning sutton1988learning ; seijen2014true . However, value-based methods struggle in continuous or large discrete space for its greedy objective. Thus, policy-based methods were introduced to model the decision-making process using a parameterized policy. As one of its best-known algorithms, Proximal Policy Optimization (PPO) schulman2017proximal is widely used in various domains, including Natural Language Processing (NLP).
Alignment for LMs has emerged as a crucial task these years, which adjusts the LMs’ generation distribution in line with human preferences. Building on the Bradley-Terry (BT) bradley1952rank Model, current alignment methods typically train models to distinguish positive and negative answers given the same instruction using RLHF/RLAIF ziegler2019fine ; ouyang2022training ; lee2023rlaif . While PPO remains the primary algorithm for alignment, its high demands for computation and memory hinders its broader use. Consequently, many improved methods have been proposed dong2023raft ; yuan2023rrhf ; zhao2023slic . Among them, DPO rafailov2024direct unifies reward modeling and policy optimization by utilizing a loss function derived from PPO, training a single model to serve as both a policy model and a reward model. Without sacrificing performance, DPO decrease the costly consumption of PPO through directly value iteration similar to a preference-based format of MC instead of TD. Nevertheless, the expensive preference-labeling process is still required. Additionally, a warm-up stage based on SFT is also necessary, which may bring trade-offs in fitting different objectives between SFT and Preference Optimization.
Improved Versions of DPO come out one after another. Efforts such as liu2023makes ; khaki2024rs ; yin2024relative ; guo2024controllable ; bansal2024comparing ; liu2024extensive try to enhance the contrastive learning by utilizing better ranking strategies, more informative data, or more number of negative samples. except for using offline data, liu2023statistical ; yuan2024self ; guo2024direct ; calandriello2024human ; chen2024self ; mitra2024orca focus on online sampling and automated label/reward collection, reducing the manual cost required for alignment. Recently, methods like ethayarajh2024kto ; hong2024reference aim to reduce DPO’s dependency on SFT warm-up by transforming its loss functions and data format into a SFT manner. These algorithms handle positive and negative samples using SFT objective and Unlikelihood Training welleck2019neural , respectively. However, the actual volume of training data is not decreased in these methods. Also, GPU-memory-consuming pair-wise data is still required, while the need for a reference model and preference-labeling for the entire answer trajectory is only eliminated in limited cases.
@@ 3.2 Preference Estimation
$\mathcal{P_{\theta}}(\rho_{0})\rightarrow\mathcal{P^{*}}(\rho_{0}):\bigg{\{}\begin{aligned} &\mathcal{P_{\theta}}(\rho_{0}):\rho_{0}\rightarrow[\pi_{\theta}(\rho_{0}),\pi_{\theta}(s_{1}^{\theta}),\pi_{\theta}(s_{2}^{\theta}),\ldots]\\ &\mathcal{P^{*}}(\rho_{0}):\rho_{0}\rightarrow[\pi^{*}(\rho_{0}),\pi^{*}(s_{1}^{*}),\pi^{*}(s_{2}^{*}),\ldots]\end{aligned}$ (4)
As the truly preferences are difficult to obtain, alignment is usually conducted based on the Preference Estimation of model and human, denoted as $\hat{\mathcal{P}_{\theta}}$ and $\mathcal{\hat{P^{*}}}$ respectively. The estimations from some common methods are listed in Table 1.
@@ 3.3 Transition Optimization
However, considering the limited data, only matrix elements representing state-action/state-state pairs contained in the dataset $\mathcal{D}$ would be aligned. Given a data sample with instruction $\rho_{0}$ and target answer with length-$N$, the objective would be:
$\forall a\in\mathcal{A},n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}:\mathcal{T}_{\theta}(a,s_{n}^{*})\rightarrow\mathcal{T^{*}}(a,s_{n}^{*})$ (8)
Or equivalent to:
$\forall n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}:T_{\theta}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ (9)
$T_{\theta}(s_{n}^{*}|\rho_{0})=\prod\limits_{i=0}^{n}\mathcal{T}_{\theta}(a_{i}^{*}|s_{i}^{*}),~{}T^{*}(s_{n}^{*}|\rho_{0})=\prod\limits_{i=0}^{n}\mathcal{T^{*}}(a_{i}^{*}|s_{i}^{*})$ (10)
@@
Consequently, the loss function can be derived from the disparities of the transition matrices between model and human. Some typical loss function are listed in Appendix A.4
Table 1: Reformulation of SFT, PPO and DPO
3.4 From SFT to Preference Optimization
@@
thus LMs tend to assign higher probabilities to its own generation than to target answer given the same initial instruction:
$\forall n\in[0,N],s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*},s_{n}^{\theta}\in\mathcal{S}_{\rho_{0}}^{\theta}:T_{\theta}(s_{n}^{*},\rho_{0})\leq T_{\theta}(s_{n}^{\theta},\rho_{0})\leq 1$ (12)
@@
Figure 2: The Training Paradigm of Different Methods. Symbol * and $\theta$ denote human and model respectively, with $a_{i}^{*}=\pi^{*}(s_{i}^{*})$ and $s_{i+1}^{*}=[s_{i}^{*},a_{i}^{*}]$, similarly for $\theta$. SFT uses priors deviating from model distribution, resulting in a more biased estimation of model preferences compared to PPO and DPO. IFT achieves a closer estimation than SFT by utilizing temporary residual connections, while maintaining the data and computational efficiency as SFT.
@@ 4 Method
While SFT is data and computation-efficient, it has an inferior approximation for both Preference Estimation and Transition Optimization. In the other side, Preference Optimization (represented by PPO and DPO) enjoys better approximation at the expense of preference data construction. We hope to make good use of their strength, using solely the target data as SFT but having similar approximation as Preference Optimization.
@@ 4.1 Intuitive Preference Estimation
A key distinction between SFT and Preference Optimization lies in whether the full distribution of model preference for each initial instruction is sampled. Contrasted to Preference Optimization, the intermediate state of the target answer used for prior in SFT may be far away from the model preference, leading to inferior outcomes.
To obtain a state estimation $\hat{s_{i}^{\theta}}$ closer to model preference, we introduce a model-based distribution disturbance function $\delta_{\theta}$ for the biased state:
@@
which can also be interpreted as a temporal residual connection. Through this approach, model can predict not only the next token from intermediate state of target answer, but also develop an intuitive sense to the entire answer generation solely based on the initial instruction, deriving more accurate Preference Estimation for model:
@@ 4.2 Dynamic Relation Propagation
With improved Preference Estimation, we achieve a Transition Optimization process closer to the original objective:
$\forall n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}:\hat{T_{\theta}}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ (19)
where $s_{0}^{*}=\rho_{0}$ and $\hat{T_{\theta}}(s_{n}^{*},\rho_{0})=\prod\limits_{i=0}^{n-1}T_{\theta}(s_{i+1}^{*},\hat{s_{i}^{\theta}})$. This objective can be optimized by the following loss function that quantifies the disparities of transition between model and human:
$\mathcal{L}(\hat{T_{\theta}};T^{*})=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{n}^{*}\sim\mathcal{S}_{\rho_{0}}^{*}}\left[\sum\limits_{n=0}^{N}\log\frac{\hat{T_{\theta}}(s_{n}^{*},\rho_{0})}{T^{*}(s_{n}^{*},\rho_{0})}\right]$ (20)
We make the same hypothesis as SFT that the optimization objective of each target intermediate state has a probability equal to 1:
$\forall n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}:T^{*}(s_{n}^{*},\rho_{0})=1=T^{*}(s_{N}^{*},\rho_{0})$ (21)
Then, the loss function can be reformulated as:
$\mathcal{L}_{\text{IFT}}(\mathcal{T}_{\theta},\delta_{\theta})=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{*}\sim\mathcal{S}_{\rho_{0}}^{*}}\left[-\sum\limits_{n=0}^{N}\sum\limits_{i=n}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))\right]$ (22)
where $a_{i}^{*}=\pi^{*}(s_{i}^{*})$. This reformulation facilitates parallel implementation, allowing IFT to achieve computational efficiency similar to SFT. See Appendix A.2 for complete derivation.
At the same time, we also prove that the objective optimized by this loss function implicitly satisfies the Bellman Equation for each state:
@@
The derivation is in Appendix A.3. This proof guarantees the optimization process closer to RLHF. Additionally, it ensures the optimized objective not only reflects the prediction accuracy of current token, but also considers the influence of the current selection on future generation, helping model acquire an intuitive understanding of generation, and a better causality for reasoning and fact-following. Additionally, a decay factor can be incorporated, as in the typical Bellman Equation, to ensure effectiveness in long trajectories.

### v3

@@ Abstract
Supervised Fine-Tuning (SFT) and Preference Optimization (PO) are key processes for aligning Language Models (LMs) with human preferences post pre-training. While SFT excels in efficiency and PO in effectiveness, they are often combined sequentially without integrating their optimization objectives. This approach ignores the opportunities to bridge their paradigm gap and take the strengths from both. In this paper, we interpret SFT and PO with two sub-processes — Preference Estimation and Transition Optimization — defined at token level within the Markov Decision Process (MDP). This modeling shows that SFT is only a special case of PO with inferior estimation and optimization. PO estimates the model’s preference by its entire generation, while SFT only scores model’s subsequent predicted tokens based on prior tokens from ground truth answer. These priors deviates from model’s distribution, hindering the preference estimation and transition optimization. Building on this view, we introduce Intuitive Fine-Tuning (IFT) to integrate SFT and PO into a single process. Through a temporal residual connection, IFT brings better estimation and optimization by capturing LMs’ intuitive sense of its entire answers. But it solely relies on a single policy and the same volume of non-preference-labeled data as SFT. Our experiments show that IFT performs comparably or even superiorly to SFT and some typical PO methods across several tasks, particularly those requires generation, reasoning, and fact-following abilities. An explainable Frozen Lake game further validates the effectiveness of IFT for getting competitive policy. Code is available at https://github.com/TsinghuaC3I/Intuitive-Fine-Tuning.
Intuitive Fine-Tuning:
Towards Simplifying Alignment into a Single Process
Ermo Hua1, Biqing Qi2,∗, Kaiyan Zhang1, Kai Tian1, Xingtai Lv1, Ning Ding1, Bowen Zhou1,2,thanks: Corresponding Author 1Department of Electronic Engineering, Tsinghua University, Beijing, China 2Shanghai AI Laboratory, Shanghai, China hem23@mails.tsinghua.edu.cn qibiqing@pjlab.org.cn, zhoubowen@tsinghua.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable powerful potential across various downstream tasks after pre-training on large-scale corpora (Brown et al., 2020; Achiam et al., 2023; Zhou and Ding, 2024). However, their instruction-following skills and trustworthiness still fall short of expectations (Bender et al., 2021; Bommasani et al., 2021; Li et al., 2022). Therefore, algorithms such as Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Ouyang et al., 2022; Lee et al., 2023) are used to further enhance LLMs’ abilities and align them better with human preferences.
@@
Figure 1: Comparison of Alignment Methods. IFT conducts alignment solely relying on positive samples and a single policy, starting from a pre-trained base model. IFT shows similar efficiency as SFT and effectiveness as PO methods.
@@
Recently, a method named Direct Preference Optimization (DPO) (Rafailov et al., 2024) was proposed to integrate Reward Modeling and Policy Optimization into one single procedure using a loss function derived from Proximal Policy Optimization (PPO) (Schulman et al., 2017). This approach demonstrates the potential to unify SFT and RLHF for the first time. Henceforth, many extended methods have been tried to realize this objective by bridging the gap between SFT and DPO. Some of them (Ethayarajh et al., 2024; Hong et al., 2024; Zhang et al., 2024) aim to transform the contrastive loss of DPO into a SFT-like cross-entropy loss, learning positive samples similar to SFT while unlearning negative samples resort to Unlikelihood Training (Welleck et al., 2019). Some others get rid of the preference-labeling process before training, switching to collect samples and labels/rewards in an online manner (Liu et al., 2023a; Yuan et al., 2024; Guo et al., 2024a; Calandriello et al., 2024; Tajwar et al., 2024), or just treating the SFT targets and online policy generations as positive and negative samples respectively (Xiong et al., 2023; Chen et al., 2024; Mitra et al., 2024; Liu et al., 2024). Nevertheless, preference-labeled pairwise data is still essential, and the need for reference model only becomes unnecessary in some cases. Thus the core differences between SFT and Preference Optimization (PO) are not eliminated thoroughly. To address this challenging issue, a deeper and more unified understanding of them are needed.
@@
In this paper, we attempt to explain the similarities and differences between SFT and PO by defining Preference Estimation and Transition Optimization in terms of state-action pairs within the Markov Decision Process (MDP) framework. Through this modeling, we demonstrate that SFT is simply a specialized case of PO with inferior estimation and optimization than other methods. To estimate the policy preference, PO collects sentence-level negative samples from policy for each initial instruction. However, SFT only samples subsequent token for each intermediate state of ground truth answer, which leads to a biased estimation of policy preference and an inferior alignment performance.
@@
Depending on this understanding, we introduce a unified alignment algorithm named Intuitive Fine-Tuning (IFT). Drawing inspiration from the human ability to grasp a intuitive sense of an answer after hearing a question, IFT employs a Temporary Residual Connection across tokens to approximate policy’s entire answer for each instruction. This approach helps IFT better estimate the policy’s preference than SFT, achieving alignment performance comparable or even superior to the sequential recipe of SFT and Preference Optimization. Additionally, IFT requires only a single policy model, and the same volume and format of data as SFT, enjoying both data and computation efficiency. These characteristics of IFT are advantageous in domains where preference data is unavailable or expensive to collect.
Our main contribution are three folds:
(1) Through defining Preference Estimation and Transition Optimization using the MDP, we demonstrate that SFT is only a special case of Preference Optimization. The similarities and differences of SFT, PPO and online/offline DPO are also compared within this framework;
(2) We introduce Intuitive Fine-tuning (IFT), a deeply unified version of SFT and Preference Optimization. It utilizes temporary residual connections to extract the model’s generation preference given the initial instructions. IFT enjoys the similar efficiency as SFT on negative sampling, but can better estimate and optimize the policy preference.
(3) Through experiments on several benchmarks, we validate that IFT performs comparably or superiorly to SFT and various Preference Optimization methods. An explainable toy-setting Frozen Lake further demonstrates the effectiveness of IFT.
Figure 2: The Training Paradigm of Different Methods. Symbol * and $\theta$ denote human and model respectively, with $a_{i}^{*}=\pi^{*}(s_{i}^{*})$ and $s_{i+1}^{*}=[s_{i}^{*},a_{i}^{*}]$, similarly for $\theta$. SFT uses priors deviating from model distribution, resulting in a more biased estimation of model preferences compared to PPO and DPO. IFT achieves a better estimation than SFT by Temporary Residual Connections across tokens. This approach passes the residual embedding from one token to the next, creating a more accurate prior while maintaining the data and computational efficiency of SFT.
@@ 2.2 Preference Estimation
$\mathcal{P_{\theta}}(\rho_{0})\rightarrow\mathcal{P^{*}}(\rho_{0})$ (4)
$\mathcal{P_{\theta}}(\rho_{0}):\rho_{0}\rightarrow[\pi_{\theta}(\rho_{0}),\pi_{\theta}(s_{1}^{\theta}),\pi_{\theta}(s_{2}^{\theta}),\ldots]$ (5)
$\mathcal{P^{*}}(\rho_{0}):\rho_{0}\rightarrow[\pi^{*}(\rho_{0}),\pi^{*}(s_{1}^{*}),\pi^{*}(s_{2}^{*}),\ldots]$
As the truly preferences are difficult to obtain, alignment is usually conducted based on the Preference Estimation of model and human, denoted as $\hat{\mathcal{P}_{\theta}}$ and $\mathcal{\hat{P^{*}}}$ respectively. The estimations from some typical methods are listed in Table 1.
@@ 2.3 Transition Optimization
However, considering the limited data, only matrix elements representing state-action/state-state pairs contained in the dataset $\mathcal{D}$ would be aligned. Given a data sample with instruction $\rho_{0}$ and target answer with length-$N$, the objective would be $\forall a\in\mathcal{A},n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}$:
$\mathcal{T}_{\theta}(a,s_{n}^{*})\rightarrow\mathcal{T^{*}}(a,s_{n}^{*})$ (9)
Or equivalent to $\forall n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}$:
$T_{\theta}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ (10)
@@
Consequently, the loss function can be derived from the disparities of the transition matrices between model and human. Some typical loss function are listed in Appendix A.4.
3 From SFT to Preference Optimization
@@
thus LMs tend to assign higher probabilities to its own generation than to target answer given the same initial instruction $\forall n\in[0,N],s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*},s_{n}^{\theta}\in\mathcal{S}_{\rho_{0}}^{\theta}$:
$T_{\theta}(s_{n}^{*},\rho_{0})\leq T_{\theta}(s_{n}^{\theta},\rho_{0})\leq 1$ (12)
@@
Table 1: Reformulation of SFT, PPO and DPO
@@ 4 Method
While SFT is data and computation-efficient, it has an inferior approximation for both Preference Estimation and Transition Optimization. On the other side, Preference Optimization (represented by PPO and DPO) enjoys better approximation at the expense of preference data construction. We hope to make good use of their strength, using solely target data as SFT but having a similar approximation as Preference Optimization. See pseudo code in Appendix B.3.
@@ 4.1 Intuitive Preference Estimation
A key distinction between SFT and Preference Optimization is whether the full distribution of model’s preference for each initial instruction is sampled. Preference Optimization samples the policy’s entire answer to estimate its preference, ensuring each generation relies on the prior adheres to the model’s distribution. But SFT only samples subsequent tokens the intermediate state of the target answer, the used prior may be far away from the model preference, leading to inferior preference estimation for model.
To obtain a prior state estimation $\hat{s_{i}^{\theta}}$ closer to model distribution, we introduce a model-based distribution disturbance function $\delta_{\theta}$ for the biased prior state:
@@
which can also be interpreted as a temporal residual connection that passes the residual embedding from one token to the next. Through this approach, model can predict not only the next token from intermediate state of target answer, but also develop an intuitive sense to the entire answer generation solely based on the initial instruction, deriving more unbiased prior and accurate Preference Estimation for model:
@@
With improved Preference Estimation, we achieve a Transition Optimization process closer to the original objective $\forall n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}$:
$\hat{T_{\theta}}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ (19)
where $s_{0}^{*}=\rho_{0}$ and $\hat{T_{\theta}}(s_{n}^{*},\rho_{0})=\prod\limits_{i=0}^{n-1}T_{\theta}(s_{i+1}^{*},\hat{s_{i}^{\theta}})$.
This objective can be optimized by the following loss function, which quantifies the disparities of transition between model and human:
$\mathcal{L}(\mathcal{T}_{\theta},\delta_{\theta})=\mathbb{E}\left[-\sum\limits_{n=0}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))\right]$ (20)
where $a_{i}^{*}=\pi^{*}(\delta^{*}(s_{i}^{*}))=\pi^{*}(s_{i}^{*})$. See Appendix A.2 for complete derivation.
@@ 4.2 Dynamic Relation Propagation
The Intuitive Preference Estimation implicitly performs Dynamic Relation Propagation, during which the generation of future tokens will be influenced by the prediction accuracy of current token.
However, limited by the parallel computing mode, the gradient map could only be built on the same time-step. Thus, the current generated tokens is unable to obtain gradient feedback from the future generated tokens. Therefore, we reformulate the loss function by a differentiable cumulative-summation to get around this limitation:
$\mathcal{L}_{\text{IFT}}=\mathbb{E}\left[-\sum\limits_{n=0}^{N}\sum\limits_{i=n}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))\right]$ (21)
This reformulation implicitly satisfies the Bellman Equation for each state, which guarantees the optimization enjoys both of the effectiveness as RLHF and efficiency as SFT:
@@
The derivation is in Appendix A.3. Additionally, a decay factor can be incorporated to ensure effectiveness in long trajectories, as in the typical Bellman Equation.
Algorithm 1 The pseudo-code of IFT
1:  Input: Initial instruction $\rho_{0}$, Ground truth $s^{*}$ with $N$ tokens: $s^{*}[1],\ldots,s^{*}[N]$
2:  Step 1: Inference One Step Ahead
3:  for $t$ in $[1,N]$ do
5:     Sample tokens: $s^{\theta}[t]=\arg\max P_{t}^{\prime}$
6:  end for
7:  Step 2: Intuitive Preference Estimation
8:  Encode $s^{*}$ and $s^{\theta}$ using Embedding Layer $E$
9:  Compute the fused embedding: $e=(1-\lambda)E(s^{*})+\lambda E(s^{\theta})$
10:  for $t$ in $[1,N]$ do
12:     Compute token loss: $\mathcal{L}_{t}=\log(P_{t}^{\prime\prime},s^{*}[t])$
13:  end for
14:  Step 3: Dynamic Relation Propagation
15:  for $t$ in $[1,N]$ do
17:  end for
18:  Output: Final loss $\mathcal{L}_{\text{IFT}}=w\cdot\mathcal{L}$
Method ARC ARC-Gen MMLU TruthfulQA WinoGrande GSM8K Avg. Mistral-7B 53.07 73.04 59.14 45.29 77.58 38.89 54.79      fine-tuning with UltraFeedback-60k + SFT 56.49 74.00 60.44 55.57 77.90 42.84 58.65 + DPO 61.86 73.54 61.02 47.98 76.64 43.89 58.28 + TDPO 56.06 73.72 60.23 43.94 77.03 41.70 55.79 + ORPO 56.66 73.98 60.57 51.77 77.19 42.30 57.70 + SimPO 59.90 73.55 52.61 47.25 78.30 37.53 55.15 + IFT 56.74 74.15 60.49 57.65 78.45 44.73 59.61 Mistral-ORPO-$\alpha$ 57.25 73.72 58.74 60.59 73.72 46.78 59.41      fine-tuning with Ultrachat-200k + UltraFeedback-60k sequentially + SFT 57.68 72.87 58.25 45.78 77.19 40.94 55.97 + SFT + SFT 58.10 72.61 58.40 48.59 76.80 43.06 56.99 + SFT + DPO 63.91 73.98 59.75 46.39 76.06 41.47 57.52 + SFT + TDPO 59.13 73.72 58.92 46.63 76.32 44.58 57.12 + SFT + ORPO 58.45 73.21 58.80 50.31 76.45 42.76 57.35 + SFT + SimPO 60.83 73.63 59.01 49.45 76.95 38.44 56.94 + SFT + IFT 58.36 73.38 58.45 52.39 78.06 43.82 58.22 Zephyr-7B-$\beta$ 67.41 72.61 58.74 53.37 74.11 33.89 57.50
Method Reference Data Alpaca-Eval Alpaca-Eval-2 TL;DR pairwise volume win-rate lc win-rate win-rate lc win-rate win-rate Mistral-7B 120k 24.72 11.57 1.25 0.35 92.03      fine-tuning with UltraFeedback-60k + SFT 120k 82.56 78.32 7.09 8.67 84.22 + DPO 120k 74.00 73.12 9.73 8.58 77.25 + TDPO 120k 65.74 51.41 4.99 3.47 70.82 + ORPO 120k 85.14 76.60 8.82 12.34 89.24 + SimPO 120k 83.08 64.30 24.47 20.31 59.13 + IFT 120k 85.18 78.78 9.95 13.27 92.63 Mistral-ORPO-$\alpha$ 120k 87.92 11.33      fine-tuning with UltraChat-200k + UltraFeedback-60k sequentially + SFT 200k 86.69 77.96 4.08 6.43 98.11 + SFT + SFT 260k 86.34 76.98 4.55 7.14 97.79 + SFT + DPO 320k 91.62 81.54 10.08 13.72 99.18 + SFT + TDPO 320k 89.80 76.44 9.25 14.15 98.89 + SFT + ORPO 320k 86.26 79.67 7.40 12.27 97.92 + SFT + SimPO 320k 88.79 68.88 19.62 23.94 98.23 + SFT + IFT 260k 88.37 81.29 10.26 14.34 98.57 Zephyr-7B-$\beta$ 320k 90.60 10.99
Table 3: Evaluation on LLM-based Benchmarks. IFT secures top two rankings in nearly all tasks, including conversation and summarization. When fine-tuned on limited data from UltraFeedback, IFT demonstrates a significant lead in TL;DR.